# The Complexity Dynamics of Grokking

### v1

@@ Abstract
We investigate the phenomenon of generalization through the lens of compression. In particular, we study the complexity dynamics of neural networks to explain grokking, where networks suddenly transition from memorizing to generalizing solutions long after over-fitting the training data. To this end we introduce a new measure of intrinsic complexity for neural networks based on the theory of Kolmogorov complexity. Tracking this metric throughout network training, we find a consistent pattern in training dynamics, consisting of a rise and fall in complexity. We demonstrate that this corresponds to memorization followed by generalization. Based on insights from rate–distortion theory and the minimum description length principle, we lay out a principled approach to lossy compression of neural networks, and connect our complexity measure to explicit generalization bounds. Based on a careful analysis of information capacity in neural networks, we propose a new regularization method which encourages networks towards low-rank representations by penalizing their spectral entropy, and find that our regularizer outperforms baselines in total compression of the dataset.
@@ 1 Introduction
A central aim of machine learning is to learn a model which explains a dataset. A good explanation—and therefore a good model—is one which generalizes to unseen data. How do we choose a good explanation? Occam’s Razor says that all else being equal, less complex (i.e. simpler) explanations are expected to generalize better. This leads us to a fundamental open question in machine learning: how do we measure the complexity of a model that has been learned to explain a dataset?
In this paper, we make this question sharp, then answer it using the tools of information theory, compression, and the Minimum Description Length principle (MDL), which we apply to neural networks to explain grokking. The MDL principle (Rissanen, 1978) is based on the following observation: any regularity in a dataset can be used to compress that dataset. By Shannon’s source coding theorem, the expected optimal coding rate for samples drawn from a distribution is equal to the entropy of the distribution. To improve the encoding, we decrease the apparent entropy of the data by finding patterns in it, which is the purpose of a model.
However, the specification of the model also requires information, so the total information is the sum of information used to specify the model, plus the remaining entropy of the dataset under the model. We have successfully compressed the data only if the total description of the model and the remaining entropy of the data is less than the naïve entropy under a random predictor. MDL tells us to minimize the sum of model complexity, denoted $C(M)$, and data entropy under the model, $H(D\mid M)$:
$\min_{M}H(D\mid M)+C(M)$ (1)
Note that if the model $M$ fully explains some subset $D_{M}\subseteq D$ then the prediction of that data becomes deterministic and $H(D_{M})=0$. Consider a trivial model which simply memorizes the data in a lookup table. Whilst this reduces the entropy of the data to zero, the model complexity is exactly equal to the original entropy. Therefore total description length has not changed, and no compression has been achieved by the model. The intuition that memorizing models are not expected to generalize is made formal via generalization bounds, which we discuss and relate to compression in section 3.
To avoid learning trivial memorizing models, practitioners introduce regularizers to the model training process – typically modified objective functions which encourage the model towards lower complexity solutions. Across a range of fields of statistical machine learning, generalization bounds of the form
@@
are ubiquitous (Vapnik, 1991; Bartlett & Mendelson, 2003). Notice the similar form of this bound and the MDL principle: if the error term is identified as the entropy of the data under the model, as is typically the case, then both MDL and the generalization bound are minimized by the sum of model complexity and data entropy under the model. That is, minimizers of the MDL principle naturally coincide with minimizers of generalization bounds.
To apply the MDL principle we require a computable measure of model complexity, $C(M)$. A widespread assumption in machine learning is that a model’s capacity is a good approximation of its complexity. However, capacity is the maximum upper-bound on model complexity. A lower bound on complexity is given by the Kolmogorov complexity. Formally, the Kolmogorov complexity $K$ of a string $s$ is equal to the length of the shortest program $p$ that prints the string:
@@
To link this back to the memorising model above, consider the string “$11111$" composed of $N$$1$"s. A program which prints this string is: print 1 $N$ times. The length of this program compared to the string’s length is asymptotically small, hence the Kolmogorov complexity of the string is small. Next consider a random string of length $N$: “$10110$", the shortest program which produces it hard-codes the string: print ‘‘$10110$". The length of this program is approximately equal to the length of the string. In fact, this is the definition of random in algorithmic information theory (AIT): a random string is one which has no description shorter than itself. Equivalently, random data is incompressible, since the shortest description of a random string is at least as long as that string.
[…]
Our central theoretical contribution is a principled approach to lossy compression of neural networks, which is best understood as a formalization of the “apparent complexity” measure introduced by Aaronson et al. (2014) using the language of algorithmic information and rate–distortion theory. We apply the resulting compression-based complexity bound to track the complexity dynamics of neural networks as they transition from memorization to generalization. The most striking example of this transition is the recently observed grokking phenomenon (Power et al., 2022), where networks suddenly transition from memorization to generalization long after over-fitting the training data.
Our main contributions are the following:
[…]
  • We explain grokking by tracking the complexity dynamics of neural networks as they transition from memorization to generalization. We find that properly regularized networks exhibit a characteristic rise and fall of complexity, while models which over-fit remain highly complex after memorizing the training data.
  • We propose a regulariztion scheme which encourages networks towards low complexity by penalizing their spectral entropy, a differentiable measure of the intrinsic dimension of the network. We compare the complexity dynamics of different regularizers, and find that our regularizer results in the least complex model amongst the baselines we study.
@@ 2 Related Work
Grokking      The grokking phenomenon (Power et al., 2022) occurs when networks suddenly transition from memorizing to generalizing solutions long after over-fitting the training data. Liu et al. (2022a) characterize conditions under which grokking occurs, and categorize learning into four distinct regimes, comprehension, grokking, memorization, and confusion by examining phase diagrams of learning performance across different hyperparameters. Liu et al. (2022b) demonstrate grokking on non-algorithmic datasets, and propose a link between a network’s parameter norm and grokking. This link is also explored by Varma et al. (2023), who explain grokking in terms of the “efficiency" of a network, as measured by the ratio of the logit to parameter norms. Nanda et al. (2023) propose task-specific loss-based “progress measures" to detect grokking, but note that we “lack a general notion of criticality that would allow us to predict when the phase transition" occurs. In this work, we suggest that complexity is the key dynamical metric to understand grokking (and generalization more broadly), and discuss why proxy measures such as the $L^{2}$ norm are insufficient.
@@
Generalization      The connection between model capacity and complexity is sometimes misunderstood amongst machine learning practitioners: the contemporary view in deep learning is that “larger models are better" (Nakkiran et al., 2019), with empirically observed scaling laws across a range of modalities and architectures (Kaplan et al., 2020; Henighan et al., 2020) which show monotonic improvement of performance with increasing model scale. This view contrasts with the classical bias-variance tradeoff, which says that “once we pass a certain threshold, larger models are worse" (Nakkiran et al., 2019). The latter view is only true assuming model capacity is equal to model complexity, and would follow from bounds of the form of Equation 2. However, this appears not to be the case. Lotfi et al. (2024) recently established the first non-vacuous generalization bounds for large language models by producing highly compressible models trained in a non-linear low-rank subspace, and bounding their generalization performance under a smoothed entropy loss by the Kolmogorov complexity. While they focus on explicit generalization bounds, we focus on the complexity dynamics of the transition from memorization to generalization to explain grokking. Goldblum et al. (2023) argue for generalization to be understood using Kolmogorov complexity, and apply the Solomonoff prior, $P(h)=2^{-K(h)}$ for hypotheses $h$ to a finite hypothesis bound (Langford & Seeger, 2001) to produce a generic generalization bound based on the Kolmogorov complexity.
@@
Coarse-graining      Aaronson et al. (2014) study the complexity dynamics of a cellular automaton which simulates the mixing of two liquids. They observe the rise and fall of complexity as entropy increases, and conjecture that this phenomenon is generic. To approximate the complexity of their automaton, they propose an “apparent complexity" measure which is the compressed file size of a coarse-grained description of the state. They informally assert that coarse-graining removes random information from the state, leaving only “interesting" information behind. We formalize their intuition using rate–distortion theory, and use the resultant complexity measure to explain grokking.
@@ 3 Complexity and Generalization
Understanding generalization is of central importance in machine learning. To what extent can we expect our models to generalize, and can we predict their performance ahead of time? Most generalization bounds are a function of model complexity, so understanding complexity is key to understanding generalization. Lotfi et al. (2024) provide one such bound, which quantifies the expected risk $R(h)$ for a given hypothesis $h$, with probability $1-\delta$:
@@
Unfortunately, naïvely compressing a model’s parameters produces poor complexity bounds, since the network contains random information left over from initialization, and from the stochastic training procedure. If we could remove the random information from a network, compressing the resulting parameters would give much tighter bounds on the network complexity. However, it is undecidable whether information is random (Li & Vitányi, 1993), so there is no generic method to remove it.
Aaronson et al. (2014) remove random information by coarse-graining their automaton’s state. They coarse-grain by smoothing and quantizing the state. The coarse-grained state is then compressed with off-the-shelf programs such as bzip2, with the compressed file size providing improved complexity estimates. They dub this metric the “apparent complexity", since it is not actually a bound on the original state’s complexity, and is arbitrary in the choice and degree of coarse-graining. We adapt their method and apply it to neural networks based on a central insight: the loss function induces a coarse-graining, which can be understood as principled lossy compression of the model.
@@ 3.1 Lossy Compression
Shannon’s theorem provides a lower-bound on coding rates which can losslessly compress signals, but more efficient coding rates are possible if some distortion is allowed. For example, JPEG compression removes high-frequency details from images which are perceptually irrelevant to the human eye, incurring some distortion. Rate–distortion theory formalizes the trade-off between coding rate, induced distortion, and perceptual quality (Blau & Michaeli, 2019), though it is typically defined over random variables. Vereshchagin & Vitányi (2004) generalize the theory of rate distortion to the algorithmic information setting. They define the algorithmic rate–distortion function $r_{x}(y)$, which is the minimum number of bits needed to specify $y$ with distortion relative to $x$ no more than ${\epsilon}$:
@@
Where $d$ is some distortion function. We will take the “input string" $x$ to be the model parameter vector $\theta$, and the “output" $y$ to be the coarse-grained parameter vector $\hat{\theta}$. The choice of distortion function $d$ is critical. In particular, we are not interested in the distortion in parameter space, $|\theta-\hat{\theta}|^{2}$. Instead, we are interested in the distortion under the loss function. Hence, we set:
@@
Substituting this distortion function into Eqn 5, we can see that the optimum can be interpreted as the least complex set of model weights $\hat{\theta}$ which achieve approximately the same performance under the loss function. We control the rate–distortion trade-off through the parameter $\epsilon$, which leads to the following distortion criterion to accept a coarse-grained approximation:
@@
$\epsilon\leftarrow\operatorname*{arg\,min}_{\epsilon}H(D\mid\hat{\theta})+r_{\theta}(\hat{\theta})$ (8)
@@ 4 Capacity and Regularization
Under this simplified model, it is clear that we can limit the information capacity of the parameters either by controlling their range through $\lambda$, or by reducing their precision through $\delta$, which can be thought of as an effective quantization level. Regularizers such as weight decay control the information capacity by limiting $\lambda$, or proxies of it such as an $L^{p}$ norm. In particular, weight decay adds an independent $L^{2}$ penalty for each parameter to the loss function. Methods which control information capacity through $\delta$ include training in reduced precision (Micikevicius et al., 2017), and quantization-aware training (Ma et al., 2024).
@@
A further complication is information which is spread across multiple parameters. Networks are typically parameterized by matrices which represent linear transformations, such that relevant information exists in their spectrum. Ultimately such information is accounted for by the total sum of parameter capacities, but it is helpful to think of information in the spectral representation.
We now introduce a regularization scheme which addresses both of these challenges.
@@ 4.1 Quantization and Noise
In contrast to methods like weight decay, which regularize the capacity of networks by controlling their range through effective penalties on $\lambda$, we introduce a simple modification to the training procedure which has previously been studied by Hinton & van Camp (1993), amongst many others: One way to set an effective lower bound on the quantization level $\delta$ from Eqn. 9 is to train networks with noisy weights. The noise prevents the parameters from forming structure below the noise threshold, limiting the information capacity “from the bottom”. To do this, we construct a set of noisy weights $\tilde{\theta}$ by adding independent Gaussian noise scaled by $\delta$ to every parameter at each forward pass:
@@ 4.2 Spectral Entropy
Low rank approximation is the problem of finding an approximating matrix $\tilde{M}$ for a matrix $M$ subject to the constraint that $\text{rank}(\tilde{M})\leq r$ for some $r$, typically $r\ll\text{rank}(M)$. Low-rank approximation of neural network weight matrices has recently been used to achieve compute– and parameter–efficient finetuning of LLMs (Hu et al., 2021; Dettmers et al., 2023), building on the insight that both the weights and update gradients live on a low-dimension subspace of the total weight space (Li et al., 2018; Aghajanyan et al., 2020).
@@
Where $\theta_{j}$ is the network parameter group of layer $j$, and $(\sigma^{j}_{i:k(\tau)},u^{j}_{i:k(\tau)},v^{j}_{i:k(\tau)})$ are the corresponding low-rank matrices of rank $k(\tau)$. Note that we can interpret the normalized singular values as a probability distribution. Then, define the spectral entropy of a matrix $H_{\text{svd}}$,
@@
Intuitively, $H_{\text{svd}}$ tells us about the effective rank of the matrix by measuring the spread of the singular values. To see that, note that if all singular values are equal, $H_{\text{svd}}$ is maximized, and if all but one of the singular values are zero, $H_{\text{svd}}$ is minimized. Roy & Vetterli (2007) formalize this by proving that $H_{\text{svd}}$ computes the effective rank of a matrix, $\text{rank}_{\text{eff}}(M)=\exp(H_{\text{svd}}(M))$, which measures the intrinsic dimension of the transformation parameterized by $M$. The spectral entropy not only provides a means of measuring the effective dimension of networks as training progresses—we can also explicitly penalize this quantity to encourage the network to learn simple, low-rank representations. In experiments, we show that our regularization method causes grokking, and produces the lowest-rank and most compressible networks of the methods we study.
@@
Input: Neural network parameters $\theta$, number of optimization steps $N$, error tolerance $\epsilon$
@@
Initialize Bayesian Optimizer $BO$;
[…]
for $i=1$ to $N$ do
[…]
@@
             if compressed_size $<$ best_compressed_size then
@@
[…]

### v2

@@ Abstract
[…]
keywords:
grokking , algorithmic complexity , phase transition
@@ 1 Introduction
When can we expect learned models to generalize well? The study of phase transitions in learning systems provides a window into fundamental questions about the nature of generalization and emergence of structure. In this work, we investigate the recently observed “grokking” phenomenon [1], where neural networks exhibit a sharp transition from memorization to generalization long after overfitting their training data. The extended delay between these two phases of learning provides an ideal setting to study the relationship between complexity and generalization in learning systems.
Simplicity has long been recognized as a common feature of good models: Occam’s Razor says that amongst models which explain some data equally well, the simplest one is expected to generalize best. Drawing inspiration from statistical physics and algorithmic information theory, we introduce a framework for measuring the intrinsic complexity of neural networks based on lossy compression and Kolmogorov complexity. Our approach parallels work by [2] which studies the complexity dynamics of physical systems. Their key insight is to measure complexity after appropriate coarse-graining of the system’s state. We formalize their insight using algorithmic rate–distortion theory, resulting in a principled lossy compression scheme for neural networks which bounds their complexity.
@@
This framework allows us to track the complexity dynamics of neural networks throughout training. We find that properly regularized networks that successfully generalize exhibit a distinct phase transition: complexity first rises as the network memorizes training data, then falls sharply as the network discovers a simpler underlying pattern that generalizes. In contrast, unregularized networks that fail to generalize remain trapped in a high-complexity memorization phase indefinitely. We establish an explicit link between our complexity measure and statistical generalization bounds, and show how these correspond to the Minimum Description Length Principle.
The Minimum Description Length (MDL) Principle [3] formalizes Occam’s Razor in the language of information theory. It says that the best model is the one which minimizes the total message length of the data. Shannon’s theorem [4] says that the expected optimal coding rate for samples drawn from a distribution is equal to the entropy of the distribution. Hence, the entropy $H$ of some data $D$ under a model $M$, $H(D|M)$ tells us the expected message length after encoding the data with the model. However, we must also consider the information cost of the model itself, which is the model complexity $C(M)$. MDL says we should take the model which minimizes the sum of these terms:
$M\leftarrow\operatorname*{arg\,min}_{M}H(D\mid M)+C(M)$ (1)
Intuitively, it is worth making our model one bit more complex if it lets us reduce the apparent entropy of the data by more than one bit, since that reduces the overall message length. Note that if the model $M$ fully explains some subset $D^{\prime}\subseteq D$ then the prediction of that data becomes deterministic and $H(D^{\prime}|M)=0$. Consider a trivial model which simply memorizes the data in a lookup table. Whilst this reduces the entropy of the data to zero, the model complexity is exactly equal to the original entropy of the data. Therefore the total description length has not changed, and no compression has been achieved by the model.
The insight that simpler models are expected to generalize better is made precise through statistical generalization bounds of the form:
$\text{test error}\leq\text{train error}+\text{model complexity}$ (2)
which are ubiquitous across a number of fields of statistical machine learning [5, 6]. When the error terms are identified with entropy, as is typically the case, these bounds reveal that models which minimize MDL are precisely those expected to generalize best. This provides the link between compression and generalization: the model which best compresses the data is the one expected to generalize best.
To apply the MDL principle, we require a measure of model complexity $C(M)$. Traditional complexity measures in machine learning such as Rademacher complexity [6] or VC dimension [7] are defined relative to specific hypothesis classes, but we seek a universal measure that applies to any computational model. The Kolmogorov complexity $K$ of a string $s$ is equal to the length of the shortest program $p$ which prints $s$ when executed:
@@
To build intuition, consider two strings of length $N$: The string ‘1111111...1’ can be generated by a short program print ‘1’ N times, so $K(s)\approx\log N$. In contrast, a random string ‘$10110$...’ has no pattern to exploit; its shortest description is itself: print ‘10110...’, giving $K(s)\approx N$. Algorithmically random strings are precisely those whose Kolmogorov complexity equals their length—they are incompressible because they contain no exploitable patterns. For neural networks, this means a network implementing a simple pattern should have low complexity despite having many parameters, while a network memorizing random labels must have high complexity. Unlike capacity measures that provide worst-case upper bounds, Kolmogorov complexity captures the actual information content of the learned function.
Figure 2: Complexity and accuracy vs steps for models regularized using our method. As models memorize, complexity increases. Complexity falls as generalization occurs. We use Algorithm 1 to coarse-grain models with $\epsilon=1$. Std. error shaded with six seeds.
To implement these ideas, we develop a principled approach to measuring neural network complexity based on lossy compression. While the Kolmogorov complexity is uncomputable, compression provides an upper bound: notice that compressed data plus its decompressor form a program that generates the original data. The compressed file-size of the network provides an upper bound on its complexity, with better compression resulting in tighter upper bounds. To compute tight compression bounds, we develop a coarse-graining procedure inspired by rate-distortion theory that removes noise while preserving functional behavior, enabling compression ratios 30-40× better than naïve approaches. This sensitivity is crucial for revealing the underlying complexity dynamics during training, shown in Fig 1. Our main contributions are:
  • 1.
    A theoretical framework connecting neural network complexity to Kolmogorov complexity through lossy compression, with explicit links to generalization bounds.
  • 2.
    Empirical demonstration of a complexity phase transition during grokking, characterized by a rise and fall pattern that distinguishes generalizing from memorizing solutions.
  • 3.
    A regularization method based on spectral entropy that encourages networks toward low-complexity representations by penalizing a differentiable measure of their effective dimension.
@@ 2 Related Work
Grokking    The grokking phenomenon [1] occurs when networks suddenly transition from memorizing to generalizing solutions long after over-fitting the training data. [8] characterize conditions under which grokking occurs, and categorize learning into four distinct regimes, comprehension, grokking, memorization, and confusion by examining phase diagrams of learning performance across different hyperparameters. [9] demonstrate grokking on non-algorithmic datasets, and propose a link between a network’s parameter norm and grokking. This link is also explored by [10], who explain grokking in terms of the “efficiency” of a network, as measured by the ratio of the logit to parameter norms. [11] propose task-specific loss-based “progress measures” to detect grokking, but note that we “lack a general notion of criticality that would allow us to predict when the phase transition” occurs. In this work, we suggest that complexity is the key dynamical metric to understand grokking (and generalization more broadly), and discuss why proxy measures such as the $L^{2}$ norm are insufficient.
@@
Generalization    The connection between model capacity and complexity is sometimes misunderstood amongst machine learning practitioners: the contemporary view in deep learning is that “larger models are better” [12], with empirically observed scaling laws across a range of modalities and architectures [13, 14] which show monotonic improvement of performance with increasing model scale. This view contrasts with the classical bias-variance tradeoff, which says that “once we pass a certain threshold, larger models are worse” [12]. The latter view is only true assuming model capacity is equal to model complexity, and would follow from bounds of the form of Equation 2. However, this appears not to be the case. [15] recently established the first non-vacuous generalization bounds for large language models by producing highly compressible models trained in a non-linear low-rank subspace, and bounding their generalization performance under a smoothed entropy loss by the Kolmogorov complexity. While they focus on explicit generalization bounds, we focus on the complexity dynamics of the transition from memorization to generalization to explain grokking. [16] argue for generalization to be understood using Kolmogorov complexity, and apply the Solomonoff prior, $P(h)=2^{-K(h)}$ for hypotheses $h$ to a finite hypothesis bound [17] to produce a generic generalization bound based on the Kolmogorov complexity.
@@
Coarse-graining    [2] study the complexity dynamics of a cellular automaton which simulates the mixing of two liquids. They observe the rise and fall of complexity as entropy increases, and conjecture that this phenomenon is generic. To approximate the complexity of their automaton, they propose an “apparent complexity” measure which is the compressed file size of a coarse-grained description of the state. They informally assert that coarse-graining removes random information from the state, leaving only “interesting” information behind. We formalize their intuition using rate–distortion theory, and use the resultant complexity measure to explain grokking.
@@ 3 Complexity and Generalization
Understanding generalization is of central importance in machine learning. To what extent can we expect our models to generalize, and can we predict their performance ahead of time? Most generalization bounds are a function of model complexity, so understanding complexity is key to understanding generalization. [15] provide one such bound, which quantifies the expected risk $R(h)$ for a given hypothesis $h$, with probability $1-\delta$:
@@
Unfortunately, naïvely compressing a model’s parameters produces poor complexity bounds, since the network contains random information left over from initialization, and from the stochastic training procedure. If we could remove the random information from a network, compressing the resulting parameters would give much tighter bounds on the network complexity. However, it is undecidable whether information is random [18], so there is no generic method to remove it.
[2] remove random information by coarse-graining their automaton’s state. They coarse-grain by smoothing and quantizing the state. The coarse-grained state is then compressed with off-the-shelf programs such as bzip2, with the compressed file size providing improved complexity estimates. They dub this metric the “apparent complexity”, since it is not actually a bound on the original state’s complexity, and is arbitrary in the choice and degree of coarse-graining. We adapt their method and apply it to neural networks based on a central insight: the loss function induces a coarse-graining, which can be understood as principled lossy compression of the model.
[…]
@@ 3.1 Lossy Compression
Shannon’s theorem provides a lower-bound on coding rates which can losslessly compress signals, but more efficient coding rates are possible if some distortion is allowed. For example, JPEG compression removes high-frequency details from images which are perceptually irrelevant to the human eye, incurring some distortion. Similarly, we want to find the least detailed model which produces acceptable performance on the training data. Since neural networks are initialized randomly and trained via stochastic optimization, much of their parameter’s information capacity stores noise. In algorithmic information theory, noise is precisely that which is incompressible, so if we naïvely compress model weights, the resultant complexity estimate from the compressed file size is dominated by the noise stored in the weights. To overcome this, we develop a lossy compression scheme which lets us formalize the trade-off between information capacity in the model and the performance on the training data as measured by the loss function. Rate–distortion theory formalizes the trade-off between coding rate, induced distortion, and perceptual quality [19], though it is typically defined over random variables. [20] generalize the theory of rate distortion to the algorithmic information setting. They define the algorithmic rate–distortion function $r_{x}(y)$, which is the minimum number of bits needed to specify $y$ with distortion relative to $x$ no more than $\epsilon$:
@@
Where $d$ is some distortion function. We will take the “input string” $x$ to be the model parameter vector $\theta$, and the “output” $y$ to be the coarse-grained parameter vector $\hat{\theta}$. The choice of distortion function $d$ is critical. In particular, we are not interested in the distortion in parameter space, $|\theta-\hat{\theta}|^{2}$. Instead, we are interested in the distortion under the loss function. Hence, we set:
@@
Substituting this distortion function into Eqn 5, we can see that the optimum can be interpreted as the least complex set of model weights $\hat{\theta}$ which induce distortion under the loss function no more than $\epsilon$. We control the rate–distortion trade-off through the parameter $\epsilon$, which leads to the following distortion criterion to accept a coarse-grained approximation:
@@
Figure 5: Complexity dynamics under varying distortion bounds $\epsilon$. As the allowed distortion increases, the estimated model complexity becomes smaller, as expected. The rise-and-fall pattern is robust under varying distortion bounds.
@@
[…]
@@ 4 Capacity and Regularization
Under this simplified model, it is clear that we can limit the information capacity of the parameters either by controlling their range through $\lambda$, or by reducing their precision through $\delta$, which can be thought of as an effective quantization level. Regularizers such as weight decay control the information capacity by limiting $\lambda$, or proxies of it such as an $L^{p}$ norm. In particular, weight decay adds an independent $L^{2}$ penalty for each parameter to the loss function. Methods which control information capacity through $\delta$ include training in reduced precision [21], and quantization-aware training [22].
@@
A further complication is information which is spread across multiple parameters. Networks are typically parameterized by matrices which represent linear transformations, such that relevant information exists in their spectrum. Ultimately such information is accounted for by the total sum of parameter capacities, but it is helpful to think of information in the spectral representation. We now introduce a regularization scheme which addresses both of these challenges.
@@ 4.1 Quantization and Noise
In contrast to methods like weight decay, which regularize the capacity of networks by controlling their range through effective penalties on $\lambda$, we introduce a simple modification to the training procedure which has previously been studied by [23], amongst many others: One way to set an effective lower bound on the quantization level $\delta$ from Eqn. 9 is to train networks with noisy weights. The noise prevents the parameters from forming structure below the noise threshold, limiting the information capacity “from the bottom”. To do this, we construct a set of noisy weights $\tilde{\theta}$ by adding independent Gaussian noise scaled by $\delta$ to every parameter at each forward pass:
@@ 4.2 Spectral Entropy
Low rank approximation is the problem of finding an approximating matrix $\tilde{M}$ for a matrix $M$ subject to the constraint that $\text{rank}(\tilde{M})\leq r$ for some $r$, typically $r\ll\text{rank}(M)$. Low-rank approximation of neural network weight matrices has recently been used to achieve compute– and parameter–efficient finetuning of LLMs [24, 25], building on the insight that both the weights and update gradients live on a low-dimension subspace of the total weight space [26, 27].
@@
[…]
@@
Intuitively, $H_{\text{svd}}$ tells us about the effective rank of the matrix by measuring the spread of the singular values. To see that, note that if all singular values are equal, $H_{\text{svd}}$ is maximized, and if all but one of the singular values are zero, $H_{\text{svd}}$ is minimized. Roy and Vetterli [28] formalize this by proving that $H_{\text{svd}}$ computes the effective rank of a matrix, $\text{rank}_{\text{eff}}(M)=\exp(H_{\text{svd}}(M))$, which measures the intrinsic dimension of the transformation parameterized by $M$. The spectral entropy not only provides a means of measuring the effective dimension of networks as training progresses—we can also explicitly penalize this quantity to encourage the network to learn simple, low-rank representations. In experiments, we show that our regularization method causes grokking, and produces the lowest-rank and most compressible networks of the methods we study.
@@
Input: Neural network parameters $\theta$, number of optimization steps NN, error tolerance $\epsilon$
@@
Initialize Bayesian Optimizer BOBO;
best_compressed_size $\leftarrow\infty$;
for $i=1$ to NN do
 $\tau_{i},\Delta_{i}\leftarrow BO.\textsc{SuggestParameters}()$;
@@
    if compressed_size << best_compressed_size then
@@
    $BO.\textsc{UpdateModel}(\tau_{i},\Delta_{i},\text{compressed\_size})$;